### v1

# Classification of thin Jordan schemes

Abstract.
Jordan schemes generalize association schemes in a similar way as Jordan algebras generalize the associative ones. It is well-known that association schemes of maximal rank are in one-to-one correspondence with groups (so-called thin schemes). In this paper, we classify Jordan schemes of maximal rank-to-order ratio and show that regular Jordan schemes correspond to a special class of Moufang loops, known as Ring Alternative loops.
The research of the second author was supported by GA ČR (Czech Science Foundation) grant EXPRO 20-31529X
1. Introduction
In 1959, Shah proposed [14] to use the Jordan matrix product as an algebraic tool for the analysis of block designs. This was a far-reaching generalization of the concept of an association scheme introduced in the same year by Bose and Mesner [3]. This concept was further developed and led to a flourishing research area called ”Algebraic Combinatorics” by Bannai and Ito in their famous monograph [2]. On the other hand, Shah’s idea was never noticed until Bailey published her book  [1] where she analysed the ”first” properties of the objects arising from Shah’s idea. In particular, she showed that the symmetrization of an association scheme always produces objects introduced by Shah. In 2001 Cameron called these objects Jordan schemes and posed the question whether any symmetric Jordan scheme is a symmetrization of an association scheme. The anwser was obtained in 2019 by Klin, Muzychuk and Reichard [10] who gave a negative answer to Cameron’s question and started a systematic investigation of Jordan schemes and, more generally, coherent Jordan configurations.
It is well known that all finite groups give rise to a special subclass of association schemes: the thin schemes [15]. They may be characterized as those that have a maximal rank-to-order ratio among all association schemes. In this paper, we classify all Jordan schemes that reach the maximum rank-to-order ratio. We also show that regular Jordan schemes with this property are in a one-to-one correspondence with a special class of Moufang loops.
The paper is organized as follows. Section 2 contains all necessary definitions and basic facts about coherent Jordan configurations and Jordan schemes. The third section is the central one of the paper. It is shown that the rank-to-order ratio of a JS is bounded by $3/2$ where equality happens if and only if the scheme is non-regular and thin. All such Jordan schemes are classified in Subsection 3.1. Subsection 3.2 deals with the thin regular Jordan schemes. It is proven there that there is a one-to-one correspondence between Jordan schemes of this type and Ring Alternative Moufang loops (RA-loops). In the fourth section it is shown that the Jordan schemes corresponding to RA-loops are autonomous, that is to say, they are not algebraic fusions of coherent configurations.
2. Background Material
2.1. Binary relations and matrices
Let $\Omega$ be a non-empty finite set. A binary relation on $\Omega$ is an arbitrary subset $s\subseteq\Omega\times\Omega$. We denote by $s{\omega}:=\{\beta\in\Omega\mid(\omega,\beta)\in s\}$ an $s$-neighborhood of $\omega$. A relation is called regular if $|s\omega|$ does not depend on $\omega$. In this case, the constant $|s\omega|$ is called valency of $s$ and will be denoted as $n_{s}$. If $s$ is a regular relation of valency one, then $s{\omega}$ contains a unique element which will be denoted as $s(\omega)$, i.e., $s{\omega}=\{s(\omega)\}$. Thus $s$ is a function on $\Omega$. All functions in this paper, including permutations of ${\sf Sym}(\Omega)$, will be considered as binary relations.
Every relation $s$ admits a transposed one $s^{t}=\{(\alpha,\beta)\in\Omega^{2}\mid(\beta,\alpha)\in s\}$. Following [5] we call a relation $s\subseteq\Omega\times\Omega$ thin if $|s\omega|\leq 1\geq|s^{t}\omega|$ holds for each $\omega\in\Omega$.
We use $ab$ for the relational product of $a,b\subseteq\Omega^{2}$ and $a\star b=ab\cup ba$ for their Jordan product. The diagonal relation $1_{\Delta},\Delta\subseteq\Omega$ is defined as $1_{\Delta}:=\{(\delta,\delta)\mid\delta\in\Delta\}$. The adjacency matrix of $s\subseteq\Omega^{2}$ is the square matrix, notation $\underline{s}$, whose rows and columns are labeled by elements of $\Omega$. Its $(\alpha,\beta)$-entry $\underline{s}_{\alpha,\beta}$ is one if $(\alpha,\beta)\in s$ and zero, otherwise.
The algebra of all $\Omega\times\Omega$ matrices over a field $\mathbb{F}$ will be denoted as $M_{\Omega}(\mathbb{F})$. It is always assumed that ${\mathsf{c}har}({\mathbb{F}})\neq 2$. The standard product of two matrices $A,B\in M_{\Omega}(\mathbb{F})$ will be written as $A\cdot B$ or $AB$. Their Jordan product $\frac{1}{2}(AB+BA)$ will denoted as $A\star B$. The Schur-Hadamard (entry-wise) product will be denoted as $A\circ B$. The identity matrix and all-one matrix will be denoted as $I_{\Omega}$ and $J_{\Omega}$, respectively. The linear span of any set of matrices ${\mathcal{S}}\subseteq M_{\Omega}({\mathbb{F}})$ will be denoted as $\langle{\mathcal{S}}\rangle$.
2.2. Rainbow algebras
Consider a vector subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ which is closed with respect to t and $\circ$ and contains $I_{\Omega},J_{\Omega}$. Every such subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ has a uniquely determined standard basis — the basis consisting of the minimal $\circ$-idempotents of the algebra $({\mathcal{A}},\circ)$. These are $0,1$-matrices $A_{1},...,A_{r}\in{\mathcal{A}}$ that are pairwise $\circ$-orthogonal. Each $A_{i}$ is the adjacency matrix of a unique binary relation $s_{i}\subseteq\Omega^{2}$, i.e., $A_{i}=\underline{s}_{i},i=1,...,r$. The set $S=\{s_{1},...,s_{r}\}$ forms a partition of $\Omega^{2}$. This partition satisfies the following conditions:
  1. (a)
    $1_{\Omega}$ is a union of some relations of $S$;
  2. (b)
    $S^{t}=S$, where $S^{t}:=\{s^{t}\mid s\in S\}$;
Following [8] we call a pair $\mathfrak{S}=(\Omega,S)$ a rainbow if it satisfies the above conditions. The elements of $S$ are called basic relations of the rainbow. The numbers $|\Omega|$ and $|S|$ are called the order and the rank of the rainbow; it has rank one if and only if $|\Omega|=1$. If the point set $\Omega$ has at least two elements, then $|S|\geq 2$ where the equality holds if and only if $S=\{1_{\Omega},\Omega^{2}\setminus 1_{\Omega}\}$. These rainbows are called trivial. The adjacency matrix of a rainbow ${\mathfrak{S}}=(\Omega,S)$ is the $\Omega\times\Omega$ matrix $A({\mathfrak{S}})$ defined via $A({\mathfrak{S}})_{\alpha,\beta}=S(\alpha,\beta)$ where $S(\alpha,\beta)$ is a unique basic relation containing $(\alpha,\beta)\in\Omega^{2}$.
The rainbow of maximal rank $|\Omega|^{2}$ is called discrete. Each basic relation of a discrete rainbow contains a unique ordered pair.
In what follows, we say that a subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ is a rainbow algebra if it contains $I_{\Omega},J_{\Omega}$ and is closed with respect to $^{t}$$\circ$. We say that $\mathcal{A}$} is symmetric if every $A\in{\mathcal{A}}$ symmetric; homogeneous if $\langle I_{\Omega}\rangle\circ{\mathcal{A}}\subseteq\langle I_{\Omega}\rangle$; regular if $\langle J_{\Omega}\rangle\cdot{\mathcal{A}}\subseteq\langle J_{\Omega}\rangle$.
Given a rainbow $\mathfrak{S}=(\Omega,S)$, the linear span $\mathbb{F}[S]:=\langle\underline{s}\rangle_{s\in S}\subseteq M_{\Omega}(\mathbb{F})$ is a rainbow algebra. Thus, there exists a one-to-one correspondence between rainbows and rainbow subalgebras of $M_{\Omega}(\mathbb{F})$. The statement below collects some properties of rainbow algberas.
Proposition 2.1.
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow and ${\mathcal{A}}=\langle\underline{s}\rangle_{s\in S}$ the corresponding rainbow algebra. Then
  1. (a)
    $\mathcal{A}$} is symmetric iff all basic relations of $\mathfrak{S}$ are symmetric;
  2. (b)
    $\mathcal{A}$} is homogeneous iff $1_{\Omega}\in S$;
  3. (c)
    $\mathcal{A}$} is regular iff all basic relations of $\mathfrak{S}$ are regular.
In what follows, we call a rainbow symmetric, homogeneous, or regular depending on which of the properties mentioned above it satisfies. Notice that the regularity of a rainbow always implies its homogeneity but not vice versa. A rainbow will be called thin if all its basic relations are thin.
Given two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega,S^{\prime})$, we say that $\mathfrak{S}$ is a fusion of $\mathfrak{S^{\prime}}$ (equivalently, $\mathfrak{S^{\prime}}$ is a fission of $\mathfrak{S}$) if $\mathbb{F}[S]\subseteq\mathbb{F}[S^{\prime}]$.
Two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega^{\prime},S^{\prime})$ are called (combinatorially) isomorphic if there exists a bijection $f:\Omega\rightarrow\Omega^{\prime}$ satisfying $f(S):=\{f(s)\mid s\in S\}=S^{\prime}$ where $f(s):=\{(f(\alpha),f(\beta))\mid(\alpha,\beta)\in s\}$.
2.3. Coherent Jordan algebras
A rainbow algebra is called coherent algebra (see [8]) if it is closed with respect to matrix multiplication $\cdot$. Similarly, a rainbow algebra will be called a coherent Jordan algebra (coherent J-algebra, for short) if it is closed with respect to Jordan multiplication $\star$.
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow, and let ${\mathcal{A}}:=\langle\underline{s}\rangle_{s\in S}$ be the corresponding rainbow algebra. We say that $\mathfrak{S}$ is a coherent configuration (respectively, a J-configuration) if $\mathcal{A}$} is a coherent algebra (respectively, a J-algebra). In both cases, the algebra $\mathcal{A}$} will be referred to as adjacency algebra of $\mathfrak{S}$. In what follows, we are going to use the abbreviations CC and JC for coherent configurations and coherent J-configurations, respectively.
If $\mathfrak{S}=(\Omega,S)$ is a coherent configuration, then for any pair $s,r\in S$ we have $\underline{s}\cdot\underline{r}=\sum_{t}p_{s,r}^{t}\underline{t}$ for suitable nonnegative integers $p_{s,r}^{t}$. These numbers are called structures constants of $\mathfrak{S}$. Due to their combinatorial meaning $p_{s,r}^{t}=|\alpha s\cap\beta r^{t}|,(\alpha,\beta)\in t$, they are also called the intersection numbers of $\mathfrak{S}$.
Similarly, one can define the intersection numbers for coherent J-configuration: $p_{\{s,r\}}^{t}=|\alpha s\cap\beta r^{t}|+|\alpha r\cap\beta s^{t}|,(\alpha,\beta)\in t$. Notice that in this case we write them as $p_{\{s,r\}}^{t}$ stressing the fact that $\star$ is commutative. They also appear as the structure constants for the Jordan product $\underline{s}\star\underline{r}=\frac{1}{2}\sum_{t}p_{\{s,r\}}^{t}\underline{t}$.
Recall that a homogeneous coherent configuration is called an association scheme (see [15]). Following this pattern, we will call a homogeneous coherent J-configuration a Jordan scheme. While it is well known that association schemes are always regular, Jordan schemes do not possess this property in general. It was shown in [12] that a symmetric JS is always regular. Regarding non-symmetric Jordan schemes, they are not necessarily regular, but have a very special structure — see [12, Proposition 2.2]. Examples of non-regular Jordan schemes also appear in Subsection 3.1 of this paper.
2.4. Coherent J-closure
It follows immediately from the definition of a coherent J-algebra that the intersection of any set of coherent J-algebras defined on $\Omega$ is again a coherent J-algebra on $\Omega$. This observation leads to the notion of the coherent J-closure $J(\mathcal{S})$ of a set $\mathcal{S}\subseteq M_{\Omega}(\mathbb{F})$ of matrices as the intersection of all coherent J-algebras containing $\mathcal{S}$.
The coherent J-closure generalizes the well-known concept of the coherent closure of $\mathcal{S}$ which is the intersection of all coherent algebras containing $\mathcal{S}$. It is often called the WL-closure, since Weisfeiler and Leman were the authors of the algorithm which computes it efficiently. For this reason, we are going to denote it as $\operatorname{WL}(\mathcal{S})$. Notice that we always have $J(\mathcal{S})\subseteq\operatorname{WL}(\mathcal{S})$.
2.5. Algebraic fusions of J-configurations
Given a J-configuration ${\mathfrak{S}}=(\Omega,S)$, a permutation $\alpha\in{\sf Sym}(S)$ is called an algebraic J-automorphism of $\mathfrak{S}$} if its linear extension to $\mathbb{C}[S]$ is an automorphism of the Jordan algebra $(\mathbb{C}[S],\star)$. All algebraic J-automorphisms of $\mathfrak{S}$} form a subgroup of ${\sf Sym}(S)$ that will be denoted by $\mathsf{JAut}({\mathfrak{S}})$. Notice that the transpose permutation $\tau:a\mapsto a^{t}$ is always an algebraic J-automorphism of $(\mathbb{C}[S],\star)$. The group $\mathsf{JAut}({\mathfrak{S}})$ consists of all permutations of ${\sf Sym}(S)$ which preserve the tensor of structure constants of Jordan multiplication with respect to the standard basis. In other words, a permutation $u\mapsto u^{\prime}$ of $S$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if
$\forall_{u,v,w\in S}:\ p_{\{u^{\prime},v^{\prime}\}}^{w^{\prime}}=p_{\{u,v\}}^{w}$
The following statement collects some properties of subgroups of $\mathsf{JAut}({\mathfrak{S}})$.
Proposition 2.2.
The following properties hold.
  1. (a)
    The transpose permutation $\tau$ belongs to the center of $\mathsf{JAut}({\mathfrak{S}})$;
  2. (b)
    For any $\Phi\leq\mathsf{JAut}({\mathfrak{S}})$ and $s\in S$ the relation $\Phi(s):=\bigcup_{\phi\in\Phi}\phi(s)$ is either symmetric or anti-symmetric;
  3. (c)
    The relations $\Phi(s),s\in S$ form a JC on $\Omega$.
Proof.
Proof of part (a). One can easily check that for any pair $r,s\in S$, it holds that
$s=r^{t}\iff(\underline{s}\star\underline{r})\circ I_{\Omega}\neq 0.$
Since any $\phi\in\mathsf{JAut}({\mathfrak{S}})$ preserves $I_{\Omega}$, we conclude that
$(\underline{s}\star\underline{r})\circ I_{\Omega}\neq 0\iff(\underline{\phi(s)}\star\underline{\phi(r)})\circ I_{\Omega}\neq 0,$
hereby, implying $s=r^{t}\iff\phi(s)=\phi(r)^{t}$.
Proof of part (b). It follows from part (a) that
$(\Phi(s))^{t}=\bigcup_{\phi\in\Phi}\phi(s)^{t}=\bigcup_{\phi\in\Phi}\phi(s^{t})=\Phi(s^{t}).$
If $s^{t}=\phi(s)$ for some $\phi\in\Phi$, then $\Phi(s^{t})=\Phi(\phi(s))=\Phi(s)$ implying that $\Phi(s)$ is symmetric.
If $s^{t}\neq\phi(s)$ for every $\phi\in\Phi$, then $\psi(s^{t})\neq\phi(s)$, and, consequently, $\psi(s^{t})\cap\phi(s)=\emptyset$ for any pair of $\psi,\phi\in\Phi$. Therefore $\Phi(s^{t})\cap\Phi(s)=\bigcup_{\psi,\phi\in\Phi}(\psi(s^{t})\cap\phi(s))=\emptyset$. Hence $\Phi(s)$ is anti-symmetric in this case.
[…]
$\phi(X\star Y)=\phi(X)\star\phi(Y),\phi(X\circ Y)=\phi(X)\circ\phi(Y),\phi(X^{t})=\phi(X)^{t}.$
Since $\phi(I_{\Omega})=I_{\Omega},\phi(J_{\Omega})=J_{\Omega}$, the subspace $^{\Phi}\!\!{\mathcal{A}}=\{X\in{\mathcal{A}}\mid\forall_{\phi\in\Phi}\ \phi(X)=X\}$ is a coherent Jordan algebra. It is easy to check that the matrices $\{\underline{\Phi(s)}\}_{s\in S}$ form the standard basis of $^{\Phi}\!\!{\mathcal{A}}$. ∎
It follows from the above statement that for any $\Phi\leq\mathsf{JAut}({\mathfrak{S}})$ the pair $^{\Phi}{{\mathfrak{S}}}:=(\Omega,\{\Phi(s)\}_{s\in S})$ is also a JC which will be called an algebraic fusion of $\mathfrak{S}$}.
2.6. Constructing Jordan configurations from coherent ones
If ${\mathfrak{S}}=(\Omega,S)$ is a CC, then it is also a JC. Recall that a permutation $\phi$ of $S$ is called an algebraic automorphism (anti-automorphism) of $\mathfrak{S}$} if its linear extension to $\mathcal{A}:=\langle\{\underline{s}\}_{s\in S}\rangle$ is an automorphism (anti-automorphism) of $(\mathcal{A},\cdot)$. The group $\mathsf{AAut}({\mathfrak{S}})$ of all algebraic automorphisms of $\mathfrak{S}$} is contained in $\mathsf{JAut}({\mathfrak{S}})$. Also, $\tau\in\mathsf{JAut}({\mathfrak{S}})$. Every anti-automorphism of $\mathfrak{S}$} is a composition of $\tau$ and some $\phi\in\mathsf{AAut}({\mathfrak{S}})$. All automorphisms and anti-automorphisms of $\mathfrak{S}$} form a subgroup of $\mathsf{JAut}({\mathfrak{S}})$ which will be denoted by $\mathsf{TAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is commutative, then $\mathsf{AAut}({\mathfrak{S}})=\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is not commutative, then $\tau\not\in\mathsf{AAut}({\mathfrak{S}})$ and $\mathsf{TAut}({\mathfrak{S}}):\mathsf{AAut}({\mathfrak{S}})]=2$. It is easy to check that $\mathsf{TAut}({\mathfrak{S}})=\mathsf{AAut}({\mathfrak{S}})\times\langle\tau\rangle$. It is obvious that $\mathsf{TAut}({\mathfrak{S}})\leq\mathsf{JAut}({\mathfrak{S}})$. In general, the inclusion could be proper. The smallest example known to us has order $16$ and was found by computer search.
If ${\mathfrak{S}}=(\Omega,S)$ is a thin association scheme, then $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. In this case a permutation $\phi\in{\sf Sym}(S)$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if it satisfies $\phi(st)\in\{\phi(s)\phi(t),\phi(t)\phi(s)\}$ for all $s,t\in S$. It was shown in [13] that any such map is either automorphism or anti-automorphism [^1] The authors are grateful to R. Lipyanski who informed them about the paper [13]. Thus for thin AS we always have $\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$.
We call a JC $\mathfrak{S}=(\Omega,S)$ non-autonomous if it is an algebraic fusion of a coherent configuration $\mathfrak{C}=(\Omega,C)$ via a subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. A JC will be called autonomous if it cannot be obtained as an algebraic fusion of some CC sharing the same point set. Notice that in [12] Jordan schemes were divided into two types: proper and improper. Improper are those which are symmetrizations of ASs. The others are proper. Under this definition, any non-symmetric JS is automatically proper. Regarding symmetric Jordan schemes, there are proper ones which are non-autonomous. The smallest example known to us is of order $24$ (we know that there are no examples of order less than $16$).
Example.
Computer calculations show that $\text{AS}(24,526)$, $\text{AS}(24,597)$, $\text{AS}(24,669)$ from the catalog of small association schemes by Hanaki and Miyamoto (see [7], available in GAP [6] via the COCO2P package [11]), each has two algebraic fusions that are proper symmetric Jordan schemes.
It looks like the proper JSs constructed in [12] are also autonomous, but we did not prove this. We can claim this only for the smallest example of order $15$ constructed there.
3. Thin Jordan schemes
It is well-known that each basic graph of an association scheme $(\Omega,S)$ is regular. Therefore we always have that $|S|\leq|\Omega|$ where the equality holds if and only if each basic relation $s\in S$ is a permutation on $\Omega$. Such a scheme is called thin and it is not difficult to check that an association scheme $(\Omega,S)$ is thin if and only if $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. Since every group has a regular representation, one can consider groups as a particular case of association schemes.
Although symmetric Jordan schemes are always regular, it is no longer true for nonsymmetric ones, as the following simple example shows.
Example.
The full matrix algebra $M_{\Omega}({\mathbb{F}})$ is a discrete coherent algebra of rank $|\Omega|^{2}$. In the case of $\Omega=\{1,2\}$ the mapping $\phi:\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}\rightarrow\begin{bmatrix}d&b\\ c&a\\ \end{bmatrix}$ is an anti-automorphism. Therefore, the subspace ${\mathcal{A}}:={}^{\langle\phi\rangle}M_{\{1,2\}}({\mathbb{F}})=\left\{\begin{bmatrix}a&b\\ c&a\\ \end{bmatrix}\,\vline\,a,b,c\in{\mathbb{F}}\right\}$ is a Jordan coherent algebra. It has order two and rank three. The corresponding coherent J-configuration $(\Omega=\{1,2\},S=\{s_{0}=I_{\Omega},s_{1}=\{(1,2)\},s_{2}=s_{1}^{t}\})$ is homogeneous but not regular. It’s rank-to-order ratio $|S|/|\Omega|$ is equal to $3/2$.
The following statement shows that this is the maximal value of a rank-to-order ratio for a non-regular JS.
Proposition 3.1.
Let $(\Omega,S)$ be a non-regular JS. Then $|S|/|\Omega|\leq 3/2$ where the equality holds if and only if all basic relations are thin.
Proof.
By [12, Proposition 2.2] and its proof there exists a partition $\Omega=\Omega_{1}\cup\Omega_{0}$ with $|\Omega_{1}|=|\Omega_{0}|$ such that every regular relation is contained in $\Omega_{0}^{2}\cup\Omega_{1}^{2}$ and each non-regular one is contained either in $\Omega_{0}\times\Omega_{1}$ or in $\Omega_{1}\times\Omega_{0}$. Thus $S$ splits into a disjoint union $S=S_{r}\cup S_{0,1}\cup S_{1,0}$ where $S_{r}$ stands for all regular relations while $S_{i,1-i}:=\{s\in S\mid\subseteq\Omega_{i}\times\Omega_{1-i}\}$.
It follows from $\sum_{s\in S_{i,1-i}}|s|=|\Omega_{i}|\cdot|\Omega_{1-i}|$ and $|s|=n_{s}|\Omega_{i}|$ that $\sum_{s\in S_{i,1-i}}n_{s}=|\Omega|/2$. Since all summands in the latter sum are non-negative integers, we conclude $|S_{i,1-i}|\leq|\Omega|/2$ for $i=0,1$.
Since a basic relation $s$ is regular if and only if $s\subseteq\Omega_{0}^{2}\cup\Omega_{1}^{2}$, we can write
$|\Omega_{0}^{2}\cup\Omega_{1}^{2}|=\sum_{s\in S_{r}}|s|\iff\frac{|\Omega|^{2}}{2}=\sum_{s\in S_{r}}n_{s}|\Omega|\iff\sum_{s\in S_{r}}n_{s}=\frac{|\Omega|}{2}.$
Arguing as before we conclude that $|S_{r}|\leq|\Omega|/2$. Combining altogether we obtain
$|S|=|S_{r}|+|S_{0,1}|+|S_{1,0}|\leq\frac{3}{2}|\Omega|.$
It follows from the above arguments that an equality $|S|=\frac{3}{2}|\Omega|$ happens if and only if $n_{s}=1$ for all $s\in S$. ∎
Since the maximal value of rank-order ratio is attained by non-regular thin Jordan schemes, our first goal is to classify all of them. This is done in the next subsection.
3.1. Non-regular thin Jordan schemes
Let us start this section by describing how every commutative homogeneous coherent algebra gives rise to a non-regular homogeneous coherent Jordan algebra.
Definition 3.2.
Let $\mathcal{C}\leq M_{n}(\mathbb{F})$ be a commutative homogeneous coherent algebra. Then we define $\mathcal{J}(\mathcal{C})\subseteq M_{2n}(\mathbb{F})$ according to:
$\mathcal{J}(\mathcal{C}):=\left\{\left[\begin{smallmatrix}A&B\\[2.84526pt] C&A\end{smallmatrix}\right]\;\middle|\;A,B,C\in\mathcal{C}\right\}.$
Proposition 3.3.
For every commutative homogeneous coherent algebra $\mathcal{C}\leq M_{n}(\mathbb{F})$ we have that $\mathcal{J}(\mathcal{C})$ is a homogeneous coherent Jordan algebra.
Proof.
Clearly, $\mathcal{J}(\mathcal{C})$ is a linear subspace of $M_{2n}(\mathbb{F})$. Since $\mathcal{C}$ itself is closed under Schur-Hadamard multiplication, so is $\mathcal{J}(\mathcal{C})$. It remains to show the closure under Jordan product: to this end, let $A,B,C,D,E,F\in\mathcal{C}$. We compute
$2\begin{bmatrix}A&B\\[2.84526pt] C&A\end{bmatrix}\star\begin{bmatrix}D&E\\[2.84526pt] F&D\end{bmatrix}=\begin{bmatrix}2AD+BF+EC&2AE+2BD\\[5.69054pt] 2DC+2AF&2AD+BF+EC\end{bmatrix}\in\mathcal{J}(\mathcal{C}).\qed$
An important special case is when $\mathcal{C}$ is the adjacency algebra of a thin association scheme coming from an abelian group $G$. In this case, to simplify notation, instead $\mathcal{J}(\mathcal{A}(G,G))$ we write $\mathcal{J}(G)$. Clearly, for every abelian group $G$ we have that $\mathcal{J}(G)$ is a non-regular thin homogeneous Jordan algebra.
Proposition 3.4.
Let $\mathfrak{S}=(\Omega,\mathcal{C})$ be a non-regular thin Jordan scheme. and let $\mathcal{A}$ be its adjacency algebra. Then there exists an abelian group $G$ and a permutation matrix $P$, such that $\mathcal{A}=P^{-1}\mathcal{J}(G)P$.
Proof.
Let $\mathfrak{J}=(\Omega,\mathcal{C})$ be a non-regular thin Jordan scheme and let $\mathcal{A}$ be its adjacency algebra. Let ${\mathcal{S}}=\{A_{i}\}_{i=1}^{r}$ be the standard basis of $\mathcal{A}$. By [12, Proposition 2.2], each element of $\mathcal{S}$} is of one of the following shapes:
$\begin{bmatrix}A&O\\[2.84526pt] O&B\end{bmatrix},\quad\begin{bmatrix}O&C\\[2.84526pt] O&O\end{bmatrix},\quad\begin{bmatrix}O&O\\[2.84526pt] D&O\end{bmatrix}$,
where $A,B,C,D$ are certain matrices of order $n$, and where $O$ denotes the zero matrix. Since $\mathfrak{S}$ is thin, the matrices $A,B,C,D$ are permutation matrices.
Let
$\mathcal{G}_{1}$ $=\left\{A\;\middle|\;\exists B\,:\,\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B\end{smallmatrix}\right]\in{\mathcal{S}}\right\},$ $\mathcal{G}_{2}$ $=\left\{B\;\middle|\;\exists A\,:\,\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B\end{smallmatrix}\right]\in{\mathcal{S}}\right\},$
$\mathcal{H}$ $=\left\{C\;\middle|\;\left[\begin{smallmatrix}O&C\\[2.84526pt] O&O\end{smallmatrix}\right]\in{\mathcal{S}}\right\},$ $\mathcal{K}$ $=\left\{D\;\middle|\;\left[\begin{smallmatrix}O&O\\[2.84526pt] D&O\end{smallmatrix}\right]\in{\mathcal{S}}\right\}.$
Note that $\mathcal{K}=\mathcal{H}^{t}=\mathcal{H}^{-1}$ and that $\sum\mathcal{G}_{1}=\sum\mathcal{G}_{2}=\sum\mathcal{H}=\sum\mathcal{K}=J$. Moreover, $|\mathcal{G}_{1}|=|\mathcal{G}_{2}|=|\mathcal{H}|=|\mathcal{K}|=n$.
Next we show that:
$\forall_{A\in\mathcal{G}_{1}}\,\exists!_{B\in\mathcal{G}_{2}}\,:\,\left[\begin{smallmatrix}A&O\\[5.69054pt] O&B\end{smallmatrix}\right]\in{\mathcal{S}}.$ (1)
Indeed, suppose that $\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{1}\end{smallmatrix}\right],\left[\begin{smallmatrix}A&O\\[2.84526pt] O&B_{2}\end{smallmatrix}\right]\in{\mathcal{S}}$, but $B_{1}\neq B_{2}$. Then $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{smallmatrix}\right]\in\mathcal{A}$. But then also
$\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}-B_{2}\end{bmatrix}\circ\begin{bmatrix}A&O\\[2.84526pt] O&B_{1}\end{bmatrix}=\begin{bmatrix}O&O\\[2.84526pt] O&B_{1}\end{bmatrix}$
As $\mathcal{A}$ is closed under taking powers of matrices, we obtain that $\left[\begin{smallmatrix}O&O\\[2.84526pt] O&I\end{smallmatrix}\right]\in\mathcal{A}$. However, this is a contradiction with the assumption that $\mathfrak{J}$ is a Jordan-scheme. It follows that $B_{1}=B_{2}$.
Thus we created a bijection between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Let us denote it by $A\mapsto\varphi(A)$. Clearly, $\mathcal{G}_{1}$ is the standard basis of a thin, homogeneous coherent $J$-algebra. The same is true for $\mathcal{G}_{2}$.
Next we show that
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,\exists_{D\in\mathcal{K}}\,:\,\left[\begin{smallmatrix}A&O\\[2.84526pt] O&\varphi(A)\end{smallmatrix}\right]=\left[\begin{smallmatrix}CD&O\\[2.84526pt] O&DC\end{smallmatrix}\right].$ (2)
To see this, observe that for all $C\in\mathcal{H}$, $D\in\mathcal{K}$ we have
$2\begin{bmatrix}O&O\\[2.84526pt] D&O\end{bmatrix}\star\begin{bmatrix}O&C\\[2.84526pt] O&O\end{bmatrix}=\begin{bmatrix}CD&O\\[2.84526pt] O&DC\end{bmatrix}$.
Since $|C\mathcal{K}|=|\mathcal{K}|=|\mathcal{G}_{1}|$, there exists a $D\in\mathcal{K}$ such that $A=CD$. By (1), for this $D$ holds $\varphi(A)=DC$.
Next we show that
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,:\,\varphi(g)=C^{-1}AC.$ (3)
To this end let $A\in\mathcal{G}_{1}$. By (2), there exists $D\in\mathcal{K}$, such that $A=CD$ and $\varphi(A)=DC$. From this follows that $D=C^{-1}A$. hence $\varphi(A)=C^{-1}AC$.
Note that conjugation of $\mathcal{A}$ by a permutation matrix $P$ ($X\mapsto PXP^{-1}$) preserves $\cdot$, $\star$, $\circ$, $(-)^{t}$. In other words, the resulting algebra is a homogeneous coherent $J$-algebra $\mathcal{A}^{\prime}=P\mathcal{A}P^{-1}$, and the corresponding Jordan schemes are combinatorially isomorphic.
Next we claim that:
$\exists_{P}\,:\,\text{$P$ is a permutation matrix, and }\forall_{A\in\mathcal{G}_{1}}\,:\,P\begin{bmatrix}A&O\\[2.84526pt] O&\varphi(A)\end{bmatrix}P^{-1}=\begin{bmatrix}A&O\\[2.84526pt] O&A\end{bmatrix}$. (4)
To see this, let $C\in\mathcal{H}$ and let $P:=\left[\begin{smallmatrix}I&O\\[2.84526pt] O&C\end{smallmatrix}\right]$. Then from (3) it follows that
$P\begin{bmatrix}A&O\\[2.84526pt] O&\varphi(A)\end{bmatrix}P^{-1}=\begin{bmatrix}I&O\\[2.84526pt] O&C\end{bmatrix}\begin{bmatrix}A&O\\[2.84526pt] O&C^{-1}AC\end{bmatrix}\begin{bmatrix}I&O\\[2.84526pt] O&C^{-1}\end{bmatrix}=\begin{bmatrix}A&O\\[2.84526pt] O&A\end{bmatrix}$.
From (4), it follows that without loss of generality we may assume that $\mathcal{G}_{1}=\mathcal{G}_{2}$, and that $\varphi$ is the identity-map. From now on we make this assumption.
A immediate consequence from (3) is, that
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC=CA,$ (5)
since $A=\varphi(A)=C^{-1}AC$.
Note now that
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC\in\mathcal{H}.$ (6)
To see this we simply compute for given $A\in\mathcal{G}_{1}$ and $C\in\mathcal{H}$:
$\begin{bmatrix}A&O\\[2.84526pt] O&A\end{bmatrix}\star\begin{bmatrix}O&C\\[2.84526pt] O&O\end{bmatrix}=\frac{1}{2}\begin{bmatrix}O&AC+CA\\[2.84526pt] O&O\end{bmatrix}=\begin{bmatrix}O&AC\\[2.84526pt] O&O\end{bmatrix}$.
Next, we show that $\mathcal{G}_{1}$ is an abelian group. To this end, let $C\in\mathcal{H}$, and let $A,A^{\prime}\in\mathcal{G}_{1}$. Then
$AA^{\prime}C=A(A^{\prime}C)=(A^{\prime}C)A=A^{\prime}AC.$
Thus, $AA^{\prime}=A^{\prime}A$, whence $A\star A^{\prime}=AA^{\prime}$, and $\mathcal{G}_{1}$ is indeed an abelian group.
It remains to show that $\mathcal{H}=\mathcal{K}=\mathcal{G}_{1}$. For this, let $C=(c_{i,j})$ be the unique element of $\mathcal{H}$ with $c_{1,1}=1$ (in other words, $C\vec{e}_{1}=\vec{e}_{1}$; this exists because $\sum\mathcal{H}=J$). Let $k\in\{1,\dots,n\}$ and let $A_{k}=(a^{(k)}_{i,j})$ be the unique element of $\mathcal{G}_{1}$ with $a^{(k)}_{1,k}=1$ (in other words, $A_{k}\vec{e}_{1}=\vec{e_{k}}$; this exists because $\sum\mathcal{G}_{1}=J$).
Since $A_{k}C\colon\vec{e}_{1}\mapsto\vec{e}_{k}$, and $CA_{k}\colon\vec{e}_{1}\mapsto C\vec{e}_{k}$, and since $A_{k}C=CA_{k}$ (by Claim 7), it follows that $C=I$. But now it follows from Claim 8 that $\mathcal{H}=\{AC\mid A\in\mathcal{G}_{1}\}=\mathcal{G}_{1}$.
As $\mathcal{K}=\mathcal{H}^{-1}$, we have also $\mathcal{K}=\mathcal{G}_{1}$. By now it is clear that $\mathcal{A}=\mathcal{J}(\mathcal{G}_{1})$. ∎
3.2. Regular thin Jordan schemes
In this subsection, we assume that $(\Omega,S)$ is a regular thin JS, that is, $S$ is a set of permutations on $\Omega$. The set $S$ contains the identity permutation $1_{\Omega}$ and is closed when transposed, that is, $s^{t}\in S$ for every $s\in S$. Notice that $s^{t}$ is a permutation inverse to $s$, that is, $s^{-1}=s^{t}$.
Since $S$ is a partition of $\Omega^{2}$ into a disjoint union of permutations, for every pair $(\alpha,\beta)\in\Omega^{2}$ the basic relation $S(\alpha,\beta)\in S$ is the unique permutation that maps $\alpha$ to $\beta$. From the definition of a JS it follows that $S(\alpha,\alpha)=1_{\Omega}$ and $S(\alpha,\beta)=S(\beta,\alpha)^{-1}$. The same definition implies that the adjacency matrix $(S(\alpha,\beta))_{\alpha,\beta\in\Omega}$ is a Latin square. Pick an arbitrary point $\omega_{0}\in\Omega$ and define a binary operation $\diamond$ on $S$ as follows $a\diamond b=S(\omega_{0},a(b(\omega_{0})))$, that is, a triple $a,b,c\in S$ satisfies $a\diamond b=c$ iff $a(b(\omega_{0}))=c(\omega_{0})$.
Remark.
It follows from the definition that the product $\diamond$ depends a choice of ”base point” $\omega_{0}$. It could be that a choice of a different base point will lead to a non-isomorphic magma.
Proposition 3.5.
The magma $(S,\diamond)$ has the following properties
  1. (1)
    $1_{\Omega}$ is neutral element of $(S,\diamond)$;
  2. (2)
    $\forall_{a\in S}\ a^{t}\diamond a=1_{\Omega}$;
  3. (3)
    $\forall_{a,b\in S}\ a^{t}\diamond(a\diamond b)=b$;
  4. (4)
    $\forall_{a,b\in S}$ the equation $x\diamond a=b$ has a unique solution $x\in S$.
Proof.
(1) We have $a\diamond 1_{\Omega}=S(\omega_{0},a(1_{\Omega}(\omega_{0})))=S(\omega_{0},a(\omega_{0}))=a$. Similarly, $1_{\Omega}\diamond a=a$.
(2) This part follows from the identity $a^{t}(a(\omega_{0}))=\omega_{0}=1_{\Omega}(\omega_{0})$.
(3) $a\diamond b=c\Leftrightarrow a(b(\omega_{0}))=c(\omega_{0})\Leftrightarrow b(\omega_{0})=a^{t}(c(\omega_{0}))\Leftrightarrow a^{t}\diamond c=b$.
(4) We have $x(a(\omega_{0}))=b(\omega_{0})$. Therefore, $x=S(a(\omega_{0}),b(\omega_{0}))$. ∎
It follows from the above statement that $(S,\diamond)$ is a loop with left inverse property. It is well-known that each element $a\in S$ of a loop determines two permutations $\ell_{a}$ and $r_{a}$ on $S$, defined naturally by $\ell_{a}(x)=a\diamond x$, and $r_{a}(x)=x\diamond a$. From (3) it follows that $\ell_{a^{t}}\ell_{a}=1_{S}$, that is, $\ell_{a^{t}}=\ell_{a}^{-1}$. From (3), it follows that the left division in the loop $(S,\diamond)$ is $a^{t}\diamond b$. The right division will be denoted as $b/a$, that is, $(b/a)\diamond a=b$. Later we will show that $b/a=b\diamond a^{t}$.
Proposition 3.6.
Let $\sigma:S\rightarrow\Omega$ be the mapping defined via $\sigma(s)=s(\omega_{0})$. Then
  1. (1)
    $\sigma$ is a bijection;
  2. (2)
    $\sigma^{-1}a\sigma=\ell_{a}$, for each $a\in S$;
  3. (3)
    The triple of mappings $\sigma,\sigma,1_{S}$ is an isotopism between the Latin squares $(v/u)_{v,u\in S}$ and $(S(\alpha,\beta))_{\alpha,\beta\in\Omega}$.
Proof.
(1) If $s\neq t\in S$, then $s\cap t=\emptyset$, implying that $\sigma(s)=s(\omega_{0})\neq t(\omega_{0})=\sigma(t)$. Thus, $\sigma$ is injective. To prove surjectivity, pick an arbitrary $\omega\in\Omega$ and set $s:=S(\omega_{0},\omega)$. Then $\sigma(s)=s(\omega_{0})=\omega$.
(2) We have to show that $a\sigma=\sigma\ell_{a}$ holds for each $a\in S$. The latter equality is equivalent to the following one $\forall_{s\in S}:\ a(\sigma(s))=\sigma(\ell_{a}(s))$. Now $a(\sigma(s))=a(s(\omega_{0}))=(a\diamond s)(\omega_{0})=\sigma(\ell_{a}(s))$.
(3) We have to show that $S(\sigma(u),\sigma(v))=1_{S}(v/u)=v/u$ holds for all $u,v\in S$. Denote $w:=v/u$. Then $w\diamond u=v$ implying $w(u(\omega_{0}))=v(\omega_{0})$. Hence $w=S(\sigma(u),\sigma(v))$. ∎
The Proposition above implies that the permutations $\{\ell_{s}\}_{s\in S}$ form a Jordan scheme on $S$ which is isomorphic to the original one: $(\Omega,S)$. Therefore $\forall_{u,v\in S}\ \underline{\ell_{u}}\cdot\underline{\ell_{v}}+\underline{\ell_{v}}\cdot\underline{\ell_{u}}=\underline{\ell_{u\diamond v}}+\underline{\ell_{v\diamond u}}$, or equivalently,
$\forall_{u,v\in S}\ \ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}=\ell_{u\diamond v}\cup\ell_{v\diamond u}.$ (7)
This equality may be rephrased as
$\forall_{u,v,w\in S}\ \{u\diamond(v\diamond w),v\diamond(u\diamond w)\}=\{(u\diamond v)\diamond w,(v\diamond u)\diamond w)\}.$ (8)
As an immediate consequence of the above property we conclude that $(S,\diamond)$ has the left alternative property: $u\diamond(u\diamond v)=(u\diamond u)\diamond v$.
For a pair $u,v\in S$ we set $A_{u,v}:=\{w\in S\mid u\diamond(v\diamond w)=(u\diamond v)\diamond w\}$. Notice that $1_{\Omega}\in A_{u,v}$. It follows from (8) that $A_{u,v}=A_{v,u}$.
3.2.1. Properties of $\diamond$
We will show here that $(S,\diamond)$ is a loop. We start with the following statement.
Proposition 3.7.
Let $u,v\in S$ be an arbitrary pair. Then the following hold:
  1. (a)
    If $u\diamond v=v\diamond u$, then $\ell_{u}\ell_{v}=\ell_{v}\ell_{u}=\ell_{u\diamond v}$ and $A_{u,v}=S$;
  2. (b)
    If $u\diamond v\neq v\diamond u$, then $\ell_{u}\ell_{v}\cap\ell_{v}\ell_{u}=\emptyset$.
Proof.
Part (a) follows from (8).
Part (b). It follows from $u\diamond v\neq v\diamond u$ that $\ell_{u\diamond v}\cap\ell_{v\diamond u}=\emptyset$. Therefore $(u\diamond v)\diamond w\neq(u\diamond v)\diamond w$. Now (8) implies the claim. ∎
Proposition 3.8.
Let $u,v\in S$ and let $A:=A_{u,v}$, $\bar{A}:=S\setminus A$. Then the following are true:
  1. (a)
    $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$ (in particular, $\ell_{u\diamond v}\subseteq\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}$);
  2. (b)
    The set $A$ is invariant with respect to $\ell_{u}$ and $\ell_{v}$.
Proof.
If $u$ and $v$ commute, then we are done by Proposition 3.7(a). So, for the rest of the proof we assume that $u,v$ do not commute.
Part (a). By (7) $\ell_{u}\ell_{v}\cup\ell_{v}\ell_{v}=\ell_{u\diamond v}\cup\ell_{v\diamond u}$.
By (8) $(u\diamond v)\diamond w\in\{u\diamond(v\diamond w),v\diamond(u\diamond w)\}$. More precisely,
$(u\diamond v)\diamond w=\left\{\begin{array}[]{rl}u\diamond(v\diamond w),&w\in A,\\ v\diamond(u\diamond w),&w\not\in A\end{array}\right$.
Equivalently, $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$.
Part (b). It follows from $u^{t}\diamond(u\diamond v)=v$ and from Part (a) that $\ell_{v}\subseteq\ell_{u^{t}}\ell_{u\diamond v}\cup\ell_{u\diamond v}\ell_{u^{t}}$. Taking into account that $\ell_{u^{t}}=\ell_{u}^{-1}$ we obtain
$\ell_{v}\subseteq\ell_{u}^{-1}(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\cup(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\ell_{u}^{-1}.$
After simplifications we obtain
$\ell_{v}\subseteq\ell_{v}1_{A}\cup\ell_{u}^{-1}\ell_{v}\ell_{u}1_{\bar{A}}\cup\ell_{u}\ell_{v}1_{A}\ell_{u}^{-1}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1}.$ (9)
Let us count the intersection of $\ell_{v}$ with each summand of (9):
[…]
This implies that $(\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1})\cap\ell_{v}=\ell_{v}$ which, in turn, yields $\ell_{u}1_{\bar{A}}\ell_{u}^{-1}=1_{\bar{A}}$, or equivalently, $\ell_{u}(\bar{A})=\bar{A}$. The equality $\ell_{v}(\bar{A})=\bar{A}$ follows from $u\leftrightarrow v$ symmetry. ∎
It follows from Proposition 3.8(b) that $u\diamond A_{u,v}=A_{u,v}$, and that $v\diamond A_{u,v}=A_{u,v}$. Together with $1_{\Omega}\in A_{u,v}$ we obtain that the subloop of $(S,\diamond)$ generated by $u$ and $v$ is contained in $A_{u,v}$. In particular, $u,v\in A_{u,v}$.
Proposition 3.9.
For all $u,b\in S$ holds $(v\diamond u)^{t}=u^{t}\diamond v^{t}$.
Proof.
It follows from $u\diamond A_{u,v}=A_{u,v}$ that $u^{t}\diamond(u\diamond A_{u,v})=u^{t}\diamond A_{u,v}$. By Proposition 3.5 $u^{t}\diamond(u\diamond A_{u,v})=A_{u,v}$, Hence $u^{t}\diamond A_{u,v}=A_{u,v}$. Analogously, $v^{t}\diamond A_{u,v}=A_{u,v}$. Now it follows from $1_{\Omega}\in A_{u,v}$ that $u^{t}\diamond v^{t}\in A_{u,v}=A_{v,u}$. Therefore $(v\diamond u)\diamond(u^{t}\diamond v^{t})=v\diamond(u\diamond(u^{t}\diamond v^{t}))=v\diamond v^{t}=1_{\Omega}$. Together with $(v\diamond u)\diamond(v\diamond u)^{t}=1_{\Omega}$ we obtain $(v\diamond u)^{t}=(u^{t}\diamond v^{t})$. ∎
Proposition 3.10.
The loop $(S,\diamond)$ is alternative, that is for all $u,v\in S$ we have
[…]
Proof.
Part (L) follows from equation (8). Parts (R) and (M) follow from the inclusion $u\in A_{u,v}=A_{v,u}$. ∎
Proposition 3.11.
$(S,\diamond)$ is a Moufang loop.
Proof.
A loop is Moufang iff it is simultaneously left and right Bol loop. We start by showing that $(S,\diamond)$ is a left Bol loop. To this end we need to prove the validity of the identity
$y\diamond(z\diamond(y\diamond x))=(y\diamond(z\diamond y))\diamond x.$ (10)
This identity is equivalent to $\ell_{y}\ell_{z}\ell_{y}=\ell_{y\diamond(z\diamond y)}$. To see that the latter identity holds, we note that there exists some $w\in S$ such that $\ell_{y}\ell_{z}\ell_{y}=\ell_{w}$ (see Proposition 5.1 in the Appendix). Further we argue that since $(S,\{\ell_{s}\}_{s\in S})$ is a Jordan scheme (in particular we use that in a Jordan scheme two relations are either equal or disjoint), we have that $\ell_{w}=\ell_{y\diamond(z\diamond y)}$ if and only if $\ell_{w}(1_{S})=\ell_{y\diamond(z\diamond y)}(1_{S})$. It remains to compute:
$\ell_{w}(1_{S})=\ell_{y}(\ell_{z}(\ell_{y}(1_{S})))=y\diamond(z\diamond y)=\ell_{y\diamond(z\diamond y)}(1_{S}).$
Hence, identity (10) holds and, $(S,\diamond)$ is indeed a left Bol loop.
To see that $(S,\diamond)$ is also a right Bol loop we need to prove the validity of the identity
$((x\diamond y)\diamond z)\diamond y=x\diamond((y\diamond z)\diamond y).$ (11)
Again we note that this is equivalent to the identity $r_{y}r_{z}r_{y}=r_{(y\diamond z)\diamond y}$. From Proposition 3.9 it follows that $r_{z}=\tau\ell_{z^{t}}\tau$ where $\tau\in\mathsf{Sym}(S)$ is the transposing involution on $S$, i.e. $\tau(x)=x^{t}$. Since $(S,\diamond)$ is a left Bol loop, we have $\ell_{y^{t}}\ell_{z^{t}}\ell_{y^{t}}=\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}$ (see above). If we conjugate both sides of this identity by $\tau$, then we obtain on one hand
$\tau\ell_{y^{t}}\ell_{z^{t}}\ell_{y^{t}}\tau=(\tau\ell_{y^{t}}\tau)(\tau\ell_{z^{t}}\tau)(\tau\ell_{y^{t}}\tau)=r_{y}r_{z}r_{y},$
and on the other hand
$\tau\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}\tau=\tau\ell_{((y\diamond z)\diamond y)^{t}}\tau=r_{(y\diamond z)\diamond y},$
as desired. Thus Identity (11) holds, too, and $(S,\diamond)$ is a right Bol loop.∎
Proposition 3.12.
If the Jordan scheme $(\Omega,S)$ is thin and symmetric, then $S$ is an abelian $2$-group and $(\Omega,S)$ is a symmetric association scheme.
Proof.
Since $a^{t}=a$, each element of the loop $(S,\diamond)$ has order two. It’s well-known that a Moufang loop of exponent two is a group. Hence $(S,\diamond)$ is an elementary abelian $2$-group. ∎
3.2.2. Relation to alternative algebras
A loop $(S,\diamond)$ is ring-alternative (an RA-loop, for short [4]) if the loop ring $RS$ is an alternative ring for any commutative associative ring $R$. It is known that each RA-loop is Moufang but the converse is not true [4]. The RA-loops were characterized in [4, Theorem 1] as follows.
Theorem 3.13.
A non-associative loop $(S,\diamond)$ is an RA loop if and only if
  1. (1)
    If three elements in $S$ associate in some order, then they associate in all orders.
  2. (2)
    If three elements $u,v,w$ do not associate, then $(u\diamond v)\diamond w=u\diamond(w\diamond v)=v\diamond(u\diamond w)$.
Theorem 3.14.
Let $(S,\diamond)$ be a loop. The permutations $\ell_{a},a\in S$ form a thin Jordan scheme if and only if $(S,\diamond)$ is an RA loop.
Proof.
$\Leftarrow$” Let $(S,\diamond)$ be an RA loop. Then it is a Moufang loop [4] and, therefore, every element $a$ has an inverse $a^{-1}$ satisfying $\forall_{a,b\in S}\ a\diamond(a^{-1}\diamond b)=b$. This implies that $\ell_{a^{-1}}=\ell_{a}^{-1}=\ell_{a}^{t}$. Together with $\ell_{1_{\Omega}}=1_{S}$ we conclude that the partition $L:=\{\ell_{a}\}_{a\in S}$ of $S^{2}$ is a rainbow. If $\diamond$ is associative, then $(L,S)$ is a thin AS, and, therefore, is a Jordan scheme as well.
If $\diamond$ is non-associative, then being a Jordan scheme is equivalent to
$\forall_{a,b\in S}:\,\underline{\ell_{a}}\cdot\underline{\ell_{b}}+\underline{\ell_{b}}\cdot\underline{\ell_{a}}=\underline{\ell_{a\diamond b}}+\underline{\ell_{b\diamond a}}.$
This is equivalent to the following relation in the loop ring $RS$:
$a\diamond(b\diamond c)+b\diamond(a\diamond c)=(a\diamond b)\diamond c+(b\diamond a)\diamond c.$ (12)
If $a,b,c$ associate in alphabetic order, then by Theorem 3.13 they associate in any order and (12) follows.
If $a,b,c$ do not associate, then by Theorem 3.13 $a\diamond(b\diamond c)=(b\diamond a)\diamond c,b\diamond(a\diamond c)=(a\diamond b)\diamond c$ and (12) holds again.
$\Rightarrow$” Assume now that the permutations $L=\{\ell_{u}\}_{u\in S}$ form a Jordan scheme. Picking $\omega_{0}=1_{\Omega}$ we obtain a binary operation $\odot$ on $L$ via
$\ell_{a}\odot\ell_{b}=\ell_{c}\iff\ell_{c}(\omega_{0})=\ell_{a}(\ell_{b}(\omega_{0}))\iff c=a\diamond b.$
Thus the mapping $a\mapsto\ell_{a}$ is an isomorphism between the magmas $(L,\odot)$ and $(S,\diamond)$ and for this reason we always have $\ell_{a}\odot\ell_{b}=\ell_{a\diamond b}$.
By Proposition 3.11 $(L,\odot)$ is a Moufang loop. We have to check the two conditions of Theorem 3.13. If $\ell_{u},\ell_{v},\ell_{w}$ associate, then they generate a subgroup222here we used the property of Moufang loop that subloop generated by three associate elements is associative of $L$, and, therefore, associate in any order. Thus the first requirement of Theorem 3.13 is satisfied.
To check the second condition of the Theorem consider a non-associate triple $\ell_{u},\ell_{v},\ell_{w}$, i.e. $\ell_{u}\odot(\ell_{v}\odot\ell_{w})\neq(\ell_{u}\odot\ell_{v})\odot\ell_{w}$. Then it follows from (8) that $\ell_{u}\odot(\ell_{v}\odot\ell_{w})=(\ell_{v}\odot\ell_{u})\odot\ell_{w}$ and $\ell_{v}\odot(\ell_{u}\odot\ell_{w})=(\ell_{u}\odot\ell_{v})\odot\ell_{w}$.
It remains to show that $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Since $\ell_{u},\ell_{v},\ell_{w}$ don’t associate, the elements $\ell_{w}^{-1},\ell_{v}^{-1},\ell_{u}^{-1}$ do not associate either. Now (8) implies $\ell_{w}^{-1}\odot(\ell_{v}^{-1}\odot\ell_{u}^{-1})=(\ell_{v}^{-1}\odot\ell_{w}^{-1})\odot\ell_{u}^{-1}$. Applying -1 to both sides we obtain $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Combining this with the equality proven in the previous paragraph we conclude that $(L,\odot)$ satisfies the second condition of the Theorem 3.13. Hence $(L,\odot)$ and, therefore, $(S,\diamond)$ are RA-loops. ∎
3.2.3. RA-loops.
The RA loops were classified in [9, Theorem 1.1]. To describe them we need to introduce a certain construction.
Let $G$ be a finite group satisfying $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ and $G^{\prime}=\{e,s\}\subseteq Z(G)$ where $e$ is the identity element of $G$. Notice that according to [9] the condition $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ always implies $G^{\prime}=\{e,s\}\subseteq Z(G)$.
Define an involution $g\mapsto g^{*},g\in G$ as follows
$g^{*}=\left\{\begin{array}[]{cc}g&\mbox{if }g\in Z(G)\\ gs&\mbox{if }g\not\in Z(G)\end{array}\right.$.
On the set $G\times\{0,1\}$ we define a binary operation $\diamond$ as follows:
$(g,x)\diamond(h,y)=\left\{\begin{array}[]{cc}(gh,0),&x=y=0\\ (hg,1),&x=0,y=1\\ (gh^{*},1)&x=1,y=0\\ (g_{0}h^{*}g,0)&x=y=1,\end{array}\right$. (13)
where $g_{0}\in Z(G)$ is an arbitrary element chosen in advance. As in [9] we denote this algebraic structure as $L(G,*,g_{0})$.
Theorem 3.15.
([9, Theorem 1.1]) The algebraic structure $(L(G,*,g_{0}),\diamond)$ is a RA-loop and any RA-loop is isomorphic to $L(G,*,g_{0})$ for suitable $G$ and $g_{0}\in Z(G)$ as described earlier.
The statement below concentrates the properties of this loop which will be needed in what follows.
Proposition 3.16.
Let $L=L(G,*,g_{0})$ be as above. Then
  1. (a)
    The map $x\mapsto x^{*}$ is an anti-automorphism of $G$;
  2. (b)
    $Z(L)=Z(G)\times\{0\}$;
  3. (c)
    $Z(L)$ coincides with right and left nucleus of $L$
Proof.
Part (a). Write $g^{*}=g\alpha(g)$ where $\alpha(g)=e$ if $g\in Z(G)$ and $\alpha(g)=s$ otherwise. Then the property that $(gh)^{*}=h^{*}g^{*}$ is equivalent to $g,h]=\alpha(g)\alpha(h)\alpha(gh)^{-1}$.
It follows from the properties of $G$ that $g,h\in G$ do not commute if and only if the cosets $gZ(G),$ $hZ(G),$ $ghZ(G)$ are pairwise distinct. In this case $g,h]=s=\alpha(g)=\alpha(h)=\alpha(gh)$ and the required equality holds.
If $g,h$ commute, then they belong to an abelian normal subgroup $N:=\langle g,h,Z(G)\rangle\trianglelefteq G$. The restriction of $\alpha$ to $N$ is a homomorpism. Therefore $\alpha(gh)^{-1}\alpha(g)\alpha(h)=e=[g,h]$, hereby ending the proof of part (a).
Part (b). According to the product formula (13), for $g\in Z(G)$ and $h\in G$ we always have $gh=hg,g^{*}=g$. Therefore,
$(g,0)\diamond(h,y)=\left\{\begin{array}[]{cc}gh,&y=0\\ hg,&y=1\end{array}\right.=\left\{\begin{array}[]{cc}hg,&y=0\\ hg^{*},&y=1\\ \end{array}\right.=(h,y)\diamond(g,0)$
implying $(g,0)\in Z(L)$.
It follows from this formula that $(g,0),g\in Z(G)$ commutes with any element of $L$. If $g\not\in Z(G)$, then $(g,0)$ and $(g,1)$ do not commute. Hence, neither $(g,0)$ nor $(g,1)$ is in $Z(L)$.
Part (c). Given that $L$ is an RA-loop, it is also a Moufang loop. Thus its left nucleus $N_{\lambda}(L)$ coincides with its right nucleus $N_{\rho}(L)$. Next we are going to show that $N_{\rho}(L)=Z(L)$. To this end, for $f,g,h\in G$, $x,y\in\{0,1\}$, we define
$\mathsf{L}$ $:=((g,x)\diamond(h,y))\diamond(f,0),$
$\mathsf{R}$ $:=(g,x)\diamond((h,y)\diamond(f,0)).$
Using (13), we may compute that
$\mathsf{L}=\left\{\begin{array}[]{cc}(gh,0)\diamond(f,0),&x=y=0\\ (hg,1)\diamond(f,0),&x=0,y=1\\ (gh^{*},1)\diamond(f,0),&x=1,y=0\\ (g_{0}h^{*}g,0)\diamond(f,0)&x=y=1\end{array}\right.=\left\{\begin{array}[]{cc}(ghf,0),&x=y=0\\ (hgf^{*},1),&x=0,y=1\\ (gh^{*}f^{*},1),&x=1,y=0\\ (g_{0}h^{*}gf,0)&x=y=1,\end{array}\right$.
$\mathsf{R}=(g,x)\diamond\left\{\begin{array}[]{cc}(hf,0),&y=0\\ (hf^{*},1),&y=1\end{array}=\right.\left\{\begin{array}[]{cc}(ghf,0),&x=y=0\\ (hf^{*}g,1),&x=0,y=1\\ (g(hf)^{*},1)&x=1,y=0\\ (g_{0}(hf^{*})^{*}g,0)&x=y=1.\end{array}\right.$.
If $f\in Z(G)$ then $f^{*}=f$ which together with $(hf)^{*}=f^{*}h^{*}$ implies $\mathsf{L}=\mathsf{R}$. Therefore, $Z(G)\times\{0\}\subseteq N_{\rho}(L)$.
If $f\not\in Z(G)$ then $f^{*}=fs\not\in Z(G)$ implying $hgf^{*}\neq hf^{*}g$ for any $g\in G$ not commuting with $f^{*}$. Therefore, $\mathsf{L}\neq\mathsf{R}$ for every $g$ with $g,f]\neq e$. Hence $(f,0)\not\in N_{\rho}(L)$ for any $f\not\in Z(G)$.
It remains to show that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. Pick arbitrary $g,h\in G$ and compute $\mathsf{L}:=((g,1)\diamond(h,1))\diamond(f,1),\mathsf{R}:=(g,1)\diamond((h,1)\diamond(f,1))$. Applying (13) to computing $\diamond$-products we obtain
$\mathsf{L}=(g_{0}h^{*}g,0)\diamond(f,1)=(fg_{0}h^{*}g,1),\mathsf{R}=(g,1)\diamond(g_{0}f^{*}h,0)=(g(g_{0}f^{*}h)^{*},1)=(gh^{*}fg_{0}^{*},1).$
It follows from $g_{0}\in Z(G)$ that $g_{0}^{*}=g_{0}$ and $\mathsf{L}=\mathsf{R}$ holds iff $fh^{*}g=gh^{*}f$. If $f\not\in Z(G)$, then choosing $h=e$ and $g\not\in C_{G}(f)$ we obtain $\mathsf{L}\neq\mathsf{R}$. If $f\in Z(G)$, then we can take $g,h\in G$ with $h^{*},g]\neq e$ to ensure $\mathsf{L}\neq\mathsf{R}$. Finally, we conclude that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$.
4. Coherent closure of thin Jordan schemes
The goal of this section is to show that the JS we’d found in the previous section are autonomous. For this purpose we compute the WL-closure $\mathfrak{C}$ of a Jordan scheme $\mathfrak{J}=(S,\{L_{s}\}_{s\in S})$ obtained from an RA-loop $(S,\diamond)$, compute the group of $\mathsf{JAut}(\mathfrak{C})$ and check that no algebraic fusion of $\mathfrak{C}$ coincides with $\mathfrak{J}$.
Proposition 4.1.
Let $\Omega$ be a finite set and $G\leq{\sf Sym}(\Omega)$ a transitive subgroup. Then WL-closure of $G$ is a CC consisting of the $2$-orbits of the group $H:=C_{{\sf Sym}(\Omega)}(G)$.
Proof.
Let $\mathfrak{X}=(\Omega,C)$ be the coherent closure of $G$. Since $H$ centralizes $G$, each $h\in H$ is an automorphism of every $g\in G$. Therefore every basic relation of $\mathfrak{X}$ is a union of $2$-orbits of $H$. We have to show that each $2$-orbit of $H$ is a union of basic relations of $\mathfrak{X}$.
Pick an arbitrary pair $(\omega_{1},\omega_{2})\in\Omega^{2}$ and define $G_{\omega_{1}\rightarrow\omega_{2}}$ to be the set of all $g\in G$ which maps $\omega_{1}$ to $\omega_{2}$, that is $G_{\omega_{1}\rightarrow\omega_{2}}=\{g\in G\mid g(\omega_{1})=\omega_{2}\}=\{g\in G\mid(\omega_{1},\omega_{2})\in g\}$. Notice that $G_{\omega_{1}\rightarrow\omega_{2}}$ is non-empty and $G_{\omega_{1}\rightarrow\omega_{2}}=gG_{\omega_{1}}$ for any $g\in G_{\omega_{1}\rightarrow\omega_{2}}$.
The relation $r_{\omega_{1},\omega_{2}}:=\bigcap_{g\in G_{\omega_{1}\rightarrow\omega_{2}}}g$ is non-empty and is the union of certain elements from $C$. We claim that its is a $2$-orbit of $H$. Fix a permutation $g\in G_{\omega_{1}\rightarrow\omega_{2}}$ and write $r_{\omega_{1},\omega_{2}}=gr_{\omega_{1},\omega_{1}}$. Obviously $r_{\omega_{1},\omega_{1}}=1_{\Omega_{1}}$ where $\Omega_{1}=\mathsf{Fix}(G_{\omega_{1}})$. Clearly, $\Omega_{1}$ is $H$-invariant implying that $1_{\Omega_{1}}$ is $H$-invariant too. Notice that the action of $H$ on $1_{\Omega_{1}}$ is semi-regular (indeed, for any $\omega\in H$, if there existed $h\in H$ and $\omega^{\prime},\omega^{\prime\prime}\in H$ such that $h(\omega)=\omega$ and $h(\omega^{\prime})=\omega^{\prime\prime}$, then for any element $g\in G_{\omega\to\omega^{\prime}}$ we would have $g(\omega)=hgh^{-1}(\omega)=h(g(\omega))=\omega^{\prime\prime}$— a contradiction).
It is well-known that $|\Omega_{1}|=[N_{G}(G_{\omega_{1}}):G_{\omega_{1}}]=|C_{{\sf Sym}(\Omega)}(G)|=|H|$. Hence $H$ acts transitively in $1_{\Omega_{1}}$ and, therefore, it also acts transitively on $g1_{\Omega_{1}}=r_{\omega_{1},\omega_{2}}$, that is, $r_{\omega_{1},\omega_{2}}$ is a $2$ orbit of $H$. ∎
Proposition 4.2.
Let $(S,\diamond)$ be a loop and $G\subseteq{\sf Sym}(S)$ the subgroup generated by $\{\ell_{s}\}_{s\in S}$. Then $C_{{\sf Sym}(S)}(G)$ coincides with the subgroup $\widehat{N}_{\rho}:=\{r_{s}\}_{s\in N_{\rho}}$ of ${\sf Sym}(S)$, where $N_{\rho}$ is the right nucleus of $(S,\diamond)$.
Proof.
Recall that $N_{\rho}:=\{z\in S\mid\forall_{x,y\in S}\ x\diamond(y\diamond z)=(x\diamond y)\diamond z\}$. It is easy to see that $N_{\rho}$ contains $1_{\Omega}$ and is $\diamond$-closed. Together with $r_{y\diamond z}=r_{y}r_{z}$, we obtain that $\widehat{N}_{\rho}$ is a subgroup of ${\sf Sym}(S)$. Notice that $\widehat{N}_{\rho}$ acts semiregularly on $S$.
A permutation $\phi\in{\sf Sym}(S)$ centralizes $G$ if and only if it commutes with $\ell_{s}$ for each $s\in S$, that is, $\phi(s\diamond u)=s\diamond\phi(u)$ for all $s,u\in S$. Substituting $u=1_{\Omega}$ to this equality, we conclude that $\phi=r_{\phi(1_{\Omega})}$, i.e. $\phi(x)=x\diamond\phi(1_{\Omega})$. Now the equality $\forall_{s,u\in S}\ \phi(s\diamond u)=s\diamond\phi(u)$ reads as $\forall_{s,u\in S}\ (s\diamond u)\diamond\phi(1_{\Omega})=s\diamond(u\diamond\phi(1_{\Omega}))$ implying that $\phi(1_{\Omega})\in N_{\rho}$. To sum up, $\Phi\in C_{{\sf Sym}(S)}(G)$ implies that $\phi=r_{\phi(1_{\Omega})}$ and $\phi(1_{\Omega})\in N_{\rho}$.
On the other hand, if $s\in N_{\rho}$ then
$\forall_{u,v\in S}\ r_{s}(\ell_{u}(v))=(u\diamond v)\diamond s=u\diamond(v\diamond s)=\ell_{u}(r_{s}(v)),$
in other words $r_{s}\ell_{u}=\ell_{u}r_{s}$. ∎
4.1. Jordan algebraic automorphisms of coherent configurations
Let $\mathfrak{C}=(\Omega,C)$ be a coherent configuration, and let $\mathcal{A}$} be its adjacency algebra (considered as a Jordan algebra). Denote by $\Omega_{1},...,\Omega_{k}$ the fibers of $\mathfrak{C}$. Then the matrices $\underline{1_{\Omega_{i}}},i=1,...,k$ are the only minimal idempotents with respect to $\star$ and $\circ$. Therefore, every algebraic automorphism $\phi$ of $\mathfrak{C}$ permutes the relations $1_{\Omega_{i}},i=1,...,k$, that is, it acts on the set of indices 1,,k1,...,k. We abuse notation and write $\phi(1_{\Omega_{i}})=1_{\Omega_{\phi(i)}}$.
Proposition 4.3.
Let $\mathfrak{C}=(\Omega,C)$ be the $2$-orbit configuration of a semi-regular subgroup $H\leq{\sf Sym}(\Omega)$. Then
  1. (a)
    $\mathfrak{C}\cong\mathsf{Inv}(\Omega_{1},H)\otimes\mathfrak{D}_{k}$ where $\Omega_{1}$ is an $H$-orbit, $k$ is the number of $H$-orbits on $\Omega$, and $\mathcal{D}_{k}$ is the discrete configuration on $k]:=\{1,\dots,k\}$;
  2. (b)
    $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$.
Proof.
Proof of part (a). Let $\Omega_{1},...,\Omega_{k}$ be a complete set of $H$-orbits on $\Omega$. Clearly, $|\Omega_{1}|=...=|\Omega_{k}|=|H|$. Notice that a point set of $\mathsf{Inv}(\Omega_{1},H)\times\mathfrak{D}_{k}$ is $\Omega_{1}\times[k]$. Thus, to build a bijection $f:\Omega_{1}\times[k]\rightarrow\Omega$ let us choose arbitrary $\omega_{i}\in\Omega_{i},i=1,...,k$. Define $f(\omega,i)=h(\omega_{i})$ where $h\in H$ is a unique element satisfying $\omega=h(\omega_{1})$.
Now, we check that $f$ maps basic relations of $(\Omega_{1},H)\otimes\mathfrak{D}_{k}$ onto basic relations of $\mathfrak{C}$. Since in both configurations all basic relations have the same cardinality, namely $|H|$, it is sufficient to prove that for any quadruple of points $(\alpha,i),(\beta,j),(\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime})\in\Omega_{1}\times[k]$ the following implication holds
$r((\alpha,i),(\beta,j))=r((\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime}))\Rightarrow r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$
It follows from the first equality that $i^{\prime}=i,j^{\prime}=j$ and $\alpha^{\prime}=h(\alpha),\beta^{\prime}=h(\beta)$ for a unique $h\in H$. Let $a,b\in H$ be elements satisfying $\alpha=a(\omega_{1}),\beta=b(\omega_{1})$. Then $\alpha^{\prime}=ha(\omega_{1}),\beta^{\prime}=hb(\omega_{1})$. Now we obtain $f(\alpha,i)=a(\omega_{i}),f(\beta,j)=b(\omega_{j})$ and $f(\alpha^{\prime},i)=ha(\omega_{i}),f(\beta^{\prime},j)=hb(\omega_{j})$. The latter equality implies $(f(\alpha^{\prime},i),f(\beta^{\prime},j))=(h(f(\alpha,i)),h(f(\beta,j)))$. Therefore, $r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$.
Proof of part (b).
First, we identify $\mathfrak{C}$ with $\mathsf{Inv}(\Omega_{1},H)\times\mathfrak{D}_{k}$. Under this identification a semi-regular action of $H$ on $\Omega_{1}\times[k]$ is given via $h(\omega,i)=(h(\omega),i)$. Denoting by $F$ the set of basic relations of $\mathsf{Inv}(\Omega_{1},H)$, we can now describe the basic relations of $\mathfrak{C}$ in the following form $f]_{i,j}=\{((\alpha,i),(\beta,j))\mid(\alpha,\beta)\in f\}$ where $f\in F$. In what follows we set $C:=\{[f]_{ij}\mid f\in F,i,j\in[k]\}$.
Since $H$ acts regularly on $\Omega_{1}$, its 2-orbit scheme is thin, i.e. each $f\in F$ is a permutation on $\Omega_{1}$ and $F$ is a regular permutation subgroup of ${\sf Sym}(\Omega_{1})$. Notice that $F$ is a centralizer of $H$ in ${\sf Sym}(\Omega_{1})$ and $F\cong H$.
Let us describe certain elements of the group $\mathsf{AAut}(\mathfrak{C})$. Given $v\in F^{k}$, define a permutation $\hat{v}\in{\sf Sym}(C)$ as follows $\hat{v}:[f]_{i,j}\mapsto[v_{i}^{-1}fv_{j}]_{i,j}$. It is easy to check that $\hat{v}$ is an algebraic automorphism of $\mathfrak{C}$. Another class of algebraic automorphisms comes from the symmetric group $S_{k}$: any permutation $\psi\in S_{k}$ permutes the basic relations in a following way $f]_{i,j}\mapsto[f]_{\psi(i),\psi(j)}$. We’ll denote these permutations as $\hat{\psi}$. The third type of algebraic automorphisms arrives from ${\sf Aut}(F)$, namely: any automorphism $\psi$ of $F$ permutes the basic relations as follows $f]_{i,j}\mapsto[\psi(f)]_{i,i}$. As before, we denote this induced permutation as $\hat{\psi}$. One can check that the three types of automorphisms described above generate a group isomorphic to $F^{k}\rtimes(S_{k}\times{\sf Aut}(F))$.
Pick an arbitrary $\phi\in\mathsf{JAut}(\mathfrak{C})$. Then $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{{\varphi}(i)}}$ for a suitable ${\varphi}\in S_{k}$. Since $\hat{\varphi}\in\mathsf{AAut}(\mathfrak{C})$, $\hat{\varphi}^{-1}\phi$ is a Jordan automorphism of $\mathfrak{C}$ which fixes each $I_{\Omega_{i}},i\in[k]$. Thus in what follows we may assume that $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{i}}$. Consider the matrices $E_{i,j}:=\frac{1}{|H|}\underline{\Omega_{i}\times\Omega_{j}}$. The linear span ${\mathcal{A}}_{0}:=\langle E_{i,j}\rangle_{i,j\in[k]}$ is a two-sided ideal of ${\mathcal{A}}:=\mathbb{C}[C]$ isomorphic to $M_{k}(\mathbb{C})$. Since $E_{i,i}$ is unique non-zero solution of the equations $X\star I_{\Omega_{i}}=X,X\star X=X,X\circ X=\frac{1}{|H|}X$ and $\phi(I_{\Omega_{i}})=I_{\Omega_{i}}$, we obtain that $\phi(E_{i,i})=E_{i,i}$ for each $i\in[k]$. Given $i\neq j\in[k]$, the equations $E_{i,i}\star X=\frac{1}{2}X,E_{j,j}\star X=\frac{1}{2}X,X\circ X=\frac{1}{|H}X$ have twp non-zero solutions $E_{i,j}$ and $E_{j,i}$. Hence for any pair $i\neq j$ either $\phi(E_{i,j})=E_{j,i},\phi(E_{j,i})=E_{i,j}$ or $\phi(E_{i,j})=E_{i,j},\phi(E_{j,i})=E_{j,i}$. Let us call a pair switched in the first case and non-switched in the latter one.
Our next step is to show that all pairs $i\neq j$ are either switched or non-switched simultaneously. This trivial if $k=2$. In the case of $k\geq 3$ it is sufficient to prove that for any three distinct elements $i,j,\ell\in[k]$ all pairs $\{i,j\},\{i,\ell\},\{j,\ell\}$ are non-switched provided that one of them is such.
Assume, towards a contradiction, that $i,j$ is a non-switched pair, while $i,\ell$ is switched. Then $\phi(E_{i,j})=E_{i,j},\phi(E_{j,i})=E_{j,i}$ and $\phi(E_{i,\ell})=E_{\ell,i},\phi(E_{\ell,i})=E_{i,\ell}$. This implies
$E_{\ell,j}=2E_{i,j}\star E_{\ell,i}=2\phi(E_{i,j})\star\phi(E_{i,\ell})=2\phi(E_{i,j}\star E_{i,\ell})=0.$
A contradiction.
The above claim implies that $\phi$ acts on ${\mathcal{A}}_{0}\cong M_{k}(\mathbb{C})$ either trivially or as a matrix transposition. Combining $\phi$ with t in the second case we may always assume that $\phi$ is trivial on ${\mathcal{A}}_{0}$, i.e. $\phi(E_{i,j})=E_{i,j}$ holds for all $i,j\in[k]$. The latter equality means that for any $i,j\in[k]$ the set of relations $\{[f]_{i,j}\}_{f\in F}$ is $\phi$ invariant, i.e. there exists a permutation $\phi_{ij}\in{\sf Sym}(F)$ such that $\phi([f]_{i,j})=[\phi_{ij}(f)]_{i,j}$.
It follows from $f]_{ii}\star[h]_{ii}=[fh]_{ii}+[hf]_{ii}$ that $\phi_{ii}(fh)\in\{\phi_{ii}(f)\phi_{ii}(h),\phi_{ii}(h)\phi_{ii}(f)\}$. Thus $\phi_{ii}$ is a half-automorphism [13] of $F$. According to [13], $\phi_{ii}$ is either automorphism or anti-automorphism.
For every pair $i\neq j$ and $f,h\in F$ we have $2[f]_{i,j}\star[h]_{j,i}=[fh]_{i,i}+[hf]_{j,j}$. Applying $\phi$ to both sides we obtain
$\begin{array}[]{rcl}\phi_{ij}(f)\phi_{ji}(h)&=&\phi_{ii}(fh)\\ \phi_{ji}(h)\phi_{ij}(f)&=&\phi_{jj}(hf)\end{array}$ (14)
for all $f,h\in F$. Substituting $f=e$ into (14) we obtain
$\phi_{ij}(e)\phi_{ji}(h)=\phi_{ii}(h),\phi_{ji}(h)\phi_{ij}(e)=\phi_{jj}(h)\Rightarrow\phi_{ij}(e)^{-1}\phi_{ii}(h)\phi_{ij}(e)=\phi_{jj}(h)$
It follows from $f_{ii}(e)=e$ that $\phi_{ij}(e)=\phi_{ji}(e)^{-1}$. Substituting $h=e$ into (14) we obtain $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}$. Together with $\phi_{ji}(h)=\phi_{ij}(e)^{-1}\phi_{ii}(h)$ we obtain
$\phi_{ij}(f)\phi_{ji}(h)=\phi_{ii}(f)\phi_{ji}(e)^{-1}\phi_{ij}(e)^{-1}\phi_{ii}(h)=\phi_{ii}(f)\phi_{ii}(h).$
Now (14) implies that $\phi_{ii}(f)\phi_{ii}(h)=\phi_{ii}(fh)$, i.e. $\phi_{ii}$ is an automorphism of $F$ for every $i\in[k]$.
Since $\phi_{11}^{-1}\in{\sf Aut}(F)$, it induces an algebraic automorphism $\widehat{\phi_{11}^{-1}}\in\mathsf{AAut}(\mathfrak{C})$ which acts via $f]_{ij}\mapsto[\phi_{11}^{-1}(f)]_{ij}$. Replacing $\phi$ by $\phi\widehat{\phi_{11}^{-1}}$, we may assume that $\phi_{11}$ is trivial. This yields us $\phi_{jj}(h)=\phi_{j1}(e)^{-1}\phi_{11}(h)\phi_{1j}(e).$ Denoting $a_{j}:=\phi_{1j}(e)\in F$ we can write $\phi_{jj}(h)=h^{a_{j}}$ for each $j\in[k]$ (here $h^{a}$ stands for $a^{-1}ha$). Also $\phi_{1j}(f)=\phi_{11}(f)\phi_{j1}(e)^{-1}=fa_{j}$.
To finish the proof it is sufficient to show that $\phi=\hat{a}$ where $a:=(a_{1},...,a_{k})\in F^{k}$. The equality $\phi=\hat{a}$ is equivalent to $\phi([f]_{ij})=\hat{a}([f]_{ij})=[a_{i}^{-1}fa_{j}]_{ij}$.
It follows from (14) that $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}=a_{i}^{-1}fa_{i}\phi_{ji}(e)^{-1}$. So, it remains to show that $\phi_{ji}(e)=a_{j}^{-1}a_{i}$. In the case of $i=j$ the equality follows from $\phi_{ii}\in{\sf Aut}(F)F)$. So, we may assume that $i\neq j$. It follows from $2[e]_{ji}=[e]_{j1}\star[e]_{1i}$ that $\phi_{ji}(e)=\phi_{j1}(e)\phi_{1i}(e)=\phi_{1j}(e)^{-1}\phi_{1i}(e)=a_{j}^{-1}a_{i}$, as desired. ∎
4.2. Thin regular Jordan schemes are autonomous
In this subsection we will prove the statement in the title. To formulate it exactly we change slightly some notation for this subsection. Namely, if $(S,\diamond)$ is a loop, then we’ll write here $\hat{s}$ instead of $L_{s}$ used previously. The set of all permutations $\{\hat{s}\}_{s\in S}$ will be denoted as $\hat{S}$.
Now we are ready to formulate the main result of this subsection.
Theorem 4.4.
[…]
Proof.
Assume towards a contradiction that $\mathfrak{S}$ is an algebraic fusion of a coherent configuration $\mathfrak{C}=(S,C)$ w.r.t. to a certain subgroup $\Phi\leq\mathsf{TAut}(\mathfrak{C})$. Then each $\hat{s}=\Phi(c)$ for each $s\in S$ and every $c\in C$ contained in $\hat{s}$. In particular, this implies that each $c\in C$ is thin. By [5, Exercise 2.7.35] every such configuration is Schurian, i.e. there exists $H\leq{\sf Sym}(S)$ such that $C$ is a $2$-orbit set of $H$. Since each $2$-orbit of $H$ is thin, $H$ is semiregular. By Proposition 4.3 $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$. Hence the subgroup $\Phi_{0}:=\Phi\cap\mathsf{AAut}(\mathfrak{C})$ has index at most two in $\Phi$. Since $\Phi_{0}\leq\mathsf{AAut}(\mathfrak{C})$, the rainbow $^{\Phi_{0}}\mathfrak{C}=(S,C_{0})$ with $C_{0}:=\{\Phi_{0}(c)\}_{c\in C}$ is a coherent configuration. Thus $\hat{S}={}^{\Phi}C\sqsubseteq C_{0}\sqsubseteq C$. Since $\mathfrak{S}$ is not a co.co., the inclusion $\hat{S}\sqsubseteq C_{0}$ is proper and $\Phi:\Phi_{0}]=2$. In particular, $\Phi_{0}$ is normal in $\Phi$, which, in turn, implies that each $\Phi$-orbit on $C$ is either a $\Phi_{0}$-orbit or a disjoint union of two $\Phi_{0}$-orbits of the same size. Therefore, each $\hat{s}\in\hat{S}$ is either a basic relation of $^{\Phi_{0}}\mathfrak{C}$ or a union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of the same size. Since $^{\Phi_{0}}\mathfrak{C}$ and $^{\Phi}\mathfrak{C}$ are distinct, there exists $\hat{s}\in\hat{S}$ which splits into a disjoint union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size, say $\hat{s}=c_{1}\cup c_{2}$. It follows from $|c_{1}|=|c_{2}|=|S|/2$ that both $c_{1}$ and $c_{2}$ are non-regular basic relations of $^{\Phi_{0}}\mathfrak{C}$. Therefore, $^{\Phi_{0}}\mathfrak{C}$ is a non-homogeneous co.co., that is, each of its basic relations is non-regular. This implies that all b.relations of $^{\Phi_{0}}\mathfrak{C}$ have size $|S|/2$.
Notice, that the coherent closure $\mathfrak{T}=(S,T)$ of $\mathfrak{S}$ satisfies $S\sqsubset T\sqsubseteq C_{0}$. By Proposition 4.2 the basic relations of $T$ coincide with $2$-orbits of the right nucleous $N_{\rho}$ of the loop $(S,\diamond)$. In particular, $|t|=|N_{\rho}|$ for each $t\in T$. It follows from Proposition 3.16 that $N_{\rho}=Z(S)$, and, therefore, $|N_{\rho}|=|S|/8$. Since each $t\in T$ is a union of basic relations of $C_{0}$, a basic relation $c\in C_{0}$ has cardinality at most $|S|/8$, contrary to $|c|=|S|/2$. A contradiction. ∎
5. Appendix
Proposition 5.1.
Let $\mathcal{J}\subseteq M_{\Omega}({\mathbb{F}})$ be a Jordan subalgebra, i.e. $\mathcal{J}\star\mathcal{J}\subseteq\mathcal{J}$. Then for any three elements $a,b,c\in\mathcal{J}$ it holds that $acb+bca\in\mathcal{J}$. In particular, $aca\in\mathcal{J}$
Proof.
The inclusion $acb+bca\in\mathcal{J}$ follows from the identity $a\star(b\star c)+b\star(a\star c)-(a\star b)\star c=\tfrac{1}{2}(acb+bca)$. ∎
We do not know whether the above statement can be generalized for $4$ and more elements. More precisely, is it true that for any $n$ elements $a_{1},...,a_{n}\in\mathcal{J}$ the sum $a_{1}\cdots a_{n}+a_{n}\cdots a_{1}$ also belongs to $\mathcal{J}$?
Meanwhile, we have a much weaker statement.
Proposition 5.2.
Let $\mathcal{J}\subseteq M_{\Omega}({\mathbb{F}})$ be a $\star$-subalgebra. Then for any $n$ elements $a_{1},...,a_{n}\in\mathcal{J}$ their symmetrization $S(a_{1},...,a_{n}):=\sum_{\pi\in S_{n}}a_{\pi(1)}\cdots a_{\pi(n)}$ belongs to $\mathcal{J}$
Proof.
Induction on $n$. We know that the claim holds for $n=2$. Suppose that it holds for $n\geq 2$. Let $a_{1},\dots,a_{n+1}\in\mathcal{J}$. Then
$2S(a_{1},\dots,a_{n+1})$ $=\sum_{i=1}^{n+1}a_{i}S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})+\sum_{i=1}^{n+1}S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})a_{i}$
$=2\sum_{i=1}^{n+1}a_{i}\star S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1}).$
By induction hypothesis, for each $i\in\{1,\dots,n+1\}$ we have $S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})\in\mathcal{J}$. Thus, $S(a_{1},\dots,a_{n+1})\in\mathcal{J}$. ∎